# UniCTokens: Boosting Personalized Understanding and Generation via Unified Concept Tokens

### v2

@@ Abstract
Personalized models have demonstrated remarkable success in understanding and generating concepts provided by users. However, existing methods use separate concept tokens for understanding and generation, treating these tasks in isolation. This may result in limitations for generating images with complex prompts. For example, given the concept $\langle bo\rangle$, generating "$\langle bo\rangle$ wearing its hat" without additional textual descriptions of its hat. We call this kind of generation personalized knowledge-driven generation. To address the limitation, we present UniCTokens, a novel framework that effectively integrates personalized information into a unified vision language model (VLM) for understanding and generation. UniCTokens trains a set of unified concept tokens to leverage complementary semantics, boosting two personalized tasks. Moreover, we propose a progressive training strategy with three stages: understanding warm-up, bootstrapping generation from understanding, and deepening understanding from generation to enhance mutual benefits between both tasks. To quantitatively evaluate the unified VLM personalization, we present UnifyBench, the first benchmark for assessing concept understanding, concept generation, and knowledge-driven generation. Experimental results on UnifyBench indicate that UniCTokens shows competitive performance compared to leading methods in concept understanding, concept generation, and achieving state-of-the-art results in personalized knowledge-driven generation. Our research demonstrates that enhanced understanding improves generation, and the generation process can yield valuable insights into understanding. Our code and dataset will be released at: https://github.com/arctanxarc/UniCTokens.
Figure 1: The capability overview of UniCTokens. UniCTokens achieves personalized understanding and generation of a unified VLM using user-provided concept images and texts. This is accomplished by fine-tuning a set of unified concept tokens, which harness the mutual benefits of understanding and generation. Notably, UniCTokens supports complex personalized knowledge-driven generation, which has never been achieved by previous methods.
@@ 1 Introduction
Personalized tasks focus on understanding and generating information that users provide. Recently, there has been growing interest in personalizing understanding models, especially in transforming general-purpose models into daily life assistants [1, 2, 3]. As a result, significant effort has been devoted to improving the personalization capabilities of Large Language Models (LLMs) [4, 5, 6] and Vision-Language Models (VLMs) [7, 8, 9, 5, 10, 11, 12]. In terms of generation, with the rapid advancement of text-to-image (T2I) models [13, 14, 15, 16, 17, 18], personalization techniques can generate highly realistic and diverse images based on user-specified concepts. They mostly employ paradigms such as test-time fine-tuning [19, 20] and retrieval augmentation [21]. LLMs, VLMs, and T2I models have shown remarkable personalization performance in their respective task domains.
Despite significant advancements in personalized generation and understanding, current methods often treat these as independent tasks, failing to effectively leverage complementary semantics [22]. Personalized understanding utilizes vision-language models (VLMs) [23, 24, 25], while personalized generation primarily employ diffusion models [26, 19, 27]. This leads to a lack of a unified personalized model for users to perform both understanding and generation. Meanwhile, personalized tasks are complex and conceptually diverse in reality, as shown in Fig. 1. For instance, when training data includes only the concept $\langle bo\rangle$, diffusion models would struggle to generate suitable images of "$\langle bo\rangle$ wearing its hat", if additional textual descriptions of its hat are provided. Additionally, concepts often include subtle visual features that are critical for identifying them. Traditional understanding models typically prioritize high-level semantic information over these subtle yet critical features [28, 29]. Ignoring the mutual information of the two tasks makes current methods insufficient for fully understanding and efficiently generating concepts provided by users.
Although a recent work Yo’Chameleon [22] achieves understanding and generation upon a unified VLM [30], it still presents several challenges in personalization:
@@
  • Second, Yo’Chameleon assesses personalized understanding and generation separately, without quantifying how understanding facilitates generation, referred as knowledge-driven generation.
@@
To better assess the personalization capabilities of the unified model, we introduce a new benchmark, UnifyBench, aimed at evaluating models’ abilities in concept understanding, concept generation, and knowledge-driven generation. When we assess our UniCTokens using UnifyBench, we consistently achieve competitive or better results than all other personalization methods. Our analysis indicates that a better understanding enhances generation, while the generation process can also provide information that supports understanding, providing valuable insights for the development of general unified models. We believe that UnifyBench will serve as a strong foundation for future research in unified model personalization. To sum up, our contributions can be concluded as:
@@
  • We construct UnifyBench, a benchmark to evaluate concept understanding, concept generation, and knowledge-driven generation of a personalized unified model.
  • We conduct extensive experiments on UnifyBench. UniCTokens demonstrates competitive performance compared to leading methods in concept understanding and generation, achieving state-of-the-art results in knowledge-driven generation.
@@ 2 Related Work
Model personalization mainly involves integrating concept-related information into the outputs of the model. Recent text-to-image methods depend on recontextualizing text conditions. Text inversion [26] utilizes soft prompt tuning for special token adjustments, while Dreambooth [19] modifies the entire network weights to ensure subject fidelity. Additionally,  [27, 31, 32, 33, 34] inject concept-related information through varied training strategies. Large Language Models and Vision-Language Models have also witnessed trends in personalization.  [35] employs a patch-based LoRA for character, while [25] utilizes a dual-tower architecture for user-aware outputs. Furthermore,  [23, 24, 21] achieve personalized responses in multimodal scenarios through fine-tuning or retrieval-augmented generation, linking user information with content in images. Yo’Chameleon [22] is the first attempt at unified personalized models. However, its separate training strategy limits cross-task information sharing. We train unified concept tokens to enhance information transfer between understanding and generation tasks.
@@ 3.1 Unified VLM Personalization with Unified Concept Tokens
To model the complexity of real-world personalization, we train unified concept tokens rather than using separate tokens for understanding and generation as Yo’Chameleon [22]. Given a target concept $C$, users provide personalized inputs comprising: the name of $C$, a set of images $\{I^{i}\}_{i=1}^{n}$ (typically 3 to 10) that exclusively depict $C$, and a set of textual descriptions $\{T^{i}\}_{i=1}^{m}$ containing additional attributes of $C$ that are not visually inferable (e.g., “$C$’s favorite hat is red”).
The goal is to train unified concept tokens $M$ that can: (1) concept understanding: generate personalized textual responses $P_{\text{text}}$ related to $C$ (e.g., answering “What is $C$ doing in the photo?”), (2) concept generation: synthesize conditional images $P_{\text{img}}$ of $C$ under various prompts, and (3) knowledge-driven generation: produce personalized knowledge-driven images $P_{\text{pimg}}$ that integrate the textual information (e.g., “$C$ wearing its favorite hat”). Detailed task and metric settings can be found in the Appendix. This process can be expressed in a formula as follows:
@@ 3.2 Stage-1: Personalized Understanding Warm-up
Soft prompt tuning has proven effective in integrating new concepts for both personalized understanding and generation tasks [36, 26]. Moreover, learnable prompts are often utilized as conduits for information transfer across task [37, 38], model [39, 40], and modality [41, 42, 43]. Based on this paradigm, we represent the concept $C$ as a prompt containing learnable tokens for unified models:
@@
[…]
To enable effective cross-task information transfer in subsequent stages, warm-up training is essential. To stabilize the training process, we adopt the token initialization strategy proposed in MC-LLaVA [24]. During training, instruction tuning is employed to optimize the initial tokens. The training dataset contains three types of Visual Question Answering (VQA) samples: (1) positive recognition VQA pairs, (2) random recognition VQA pairs, and (3) conversational VQA pairs. Detailed descriptions of the data construction process can be found in MC-LLaVA [24] Notably, user-provided textual information $\{T^{i}\}_{i=1}^{m}$ is transformed by LLMs into a set of text-only QA pairs, which are also incorporated into training. Implementation details are provided in the Appendix.
@@ 3.3 Stage-2: Bootstrap Generative Learning from Understanding
Training solely on understanding tasks does not equip the unified model to directly generate images containing $C$. Existing personalized unified models [22] require a substantial number of samples ($\sim$1,000) to train a single concept, which is impractical for real-world applications. For humans, a preliminary understanding of concepts facilitates artistic creation. The more thoroughly humans comprehend a concept, the more accurately they can replicate it in their artwork. Thus, we aim to explore the potential of leveraging understanding information to facilitate generation.
Directly training generation tasks on shared tokens presents a straightforward strategy. However, this direct training approach results in the model losing its general conditioning capability on $C$ and significantly diminishes performance on understanding tasks. We posit that this discrepancy is due to variations in task distributions, which encompass specific information that cannot be directly shared. Thus, we integrate new tokens to enhance model training tailored for text-to-image (T2I). Inspired by DreamBooth [19], the prompt for Stage 2 of T2I training is formalized as follows:
[…]
@@ 3.4 Stage-3: Deepen Understanding Representation from Generation
Existing literature [28] shows that understanding models focus primarily on high-level semantic information, while the generation model is more responsive to low-level features. Experimental findings indicate that the training in Stage 1 may not effectively address tasks related to low-level characteristics, such as fine-grained recognition. After completing the generation training, model is capable of producing visually similar images, which may have the potential to enhance its understanding capabilities through effective utilization of information derived from generation process.
Show-o [46] utilizes a decode methodology following MaskGIT [47], removing masked tokens to derive the final image. To generate an image $x_{T}$, the process can be formularized as follows:
@@
[…]
@@ 4.1 Experiment Setup
Implement Details. We set the number of learnable tokens as $K=16$, $M=8$, and $N=8$ respectively. All training stages are optimized using AdamW and each stage is trained for 20 epochs. The batch size is set to 4 for understanding tasks in stage 1, and 1 for both stage 2 and stage 3, as well as for T2I generation. All experiments are conducted on A800 GPUs. For the backbone, we adopt Show-o512x512 [46] as the base model. More training details can be found in the Appendix.
Dataset. We collect Unifybench of 20 concepts: Person (10), Pets (5), Objects (5). Each concept is associated with 10~15 images for training and testing. In addition, each concept is annotated with 1~2 extra attributes that are not visually inferable from training images (e.g., “this person owns a dog”). To comprehensively evaluate both understanding and T2I capabilities, We design specific test data for each task. This includes standard understanding QA pairs and generation prompts, as well as personalized knowledge-driven queries. Please refer to Appendix for more details on our dataset.
Baselines. Our approach is primarily evaluated against four distinct categories of methods. The most direct comparison utilizes a unified base model incorporating personalized text and image prompts. Show-o’s inability to support interleaved image prompts precluded this comparison. The textual prompts are derived from the captions generated by GPT-4o for each concept and subsequently summarized to meet the required token length. Another significant baseline for comparison is the recent unified customized model, Yo’chameleon [22]. We retrain the model according to the original paper, utilizing 1,000 images per concept and a 7B base model. Additionally, we evaluate the current GPT-4o on our benchmark to serve as an upper bound. Lastly, we conduct comparisons between models focused solely on understanding or generation. More details are provided in the Appendix.
Metrics. For personalized knowledge-driven image generation, we use VLMs to score the generated images (from 0 to 1) based on their consistency with the extra attributes of concepts embedded in the prompt. Metrics for separate personalized understanding and generation are detailed in Appendix. Final results are obtained by averaging the scores across all concepts for each evaluation metric.
@@ 4.2 Our Unified Personalized Benchmark (UnifyBench)
Type Methods Model Size Token Training Images Personalized Understanding Personalized Generation Rec. VQA QA Pure Gen. People Gen. T2I Weight BLEU GPT BLEU GPT CLIP-I CLIP-T DINO Face-Simi Score CLIP-I Upper Bound GPT-4o+TP 200B $\sim$100 - 0.742 0.473 0.676 0.610 0.685 0.689 0.301 0.626 0.198 0.780 0.690 GPT-4o+IP 200B $\sim$1,000 - 0.773 0.543 0.685 0.589 0.652 0.794 0.310 0.722 0.559 0.782 0.792 Real Images - - - - - - - - 0.832 - 0.728 0.739 - 0.832 Und. Only Yo’LLaVA 13B 16 $\sim$100 0.919 0.609 0.629 0.612 0.593 - - - - - - MC-LLaVA 13B 16 $\sim$10 0.924 0.628 0.637 0.601 0.583 - - - - - - RAP-MLLM 13B $\sim$1,000 - 0.940 0.616 0.616 0.712 0.722 - - - - - - Qwen2.5-VL + TP 3B $\sim$100 - 0.660 0.407 0.727 0.574 0.774 - - - - - - Yo’LLaVA(Phi-1.5) 1.3B 16 $\sim$ 100 0.765 0.488 0.497 0.510 0.494 - - - - - - Gen.  Only Text inversion 1.0B - $\sim$10 - - - - - 0.630 0.247 0.569 0.371 0.070 0.628 DreamBooth (SD) 1.0B - $\sim$10 - - - - - 0.649 0.281 0.591 0.436 0.071 0.650 Unified Model Chamaleon+TP 7B $\sim$100 - 0.690 0.413 0.488 0.509 0.564 0.547 0.176 0.509 0.011 0.329 0.549 Chameleon+IP 7B $\sim$1,000 - 0.493 0.445 0.498 0.407 0.535 0.523 0.160 0.469 0.066 0.299 0.499 Show-o+TP 1.3B $\sim$100 - 0.566 0.461 0.409 0.504 0.579 0.664 0.264 0.553 0.048 0.770 0.660 Yo’Chameleon 7B 32 $\sim$1,000 0.764 0.474 0.507 0.510 0.581 0.697 0.236 0.590 0.224 0.266 0.698 Ours 1.3B 32 $\sim$10 0.790 0.503 0.523 0.544 0.603 0.750 0.282 0.646 0.334 0.359 0.749
Table 1: Quantitative Comparison on UnifyBench. TP = Text Prompt. IP = Image Prompt. The best and second best performances are highlighted.
@@
This task requires responding to queries with concept identifiers and images. Following [23, 36, 24], we conduct experiments on personalized recognition, VQA, and QA tasks. As demonstrated in Tab. 1, our proposed UniCTokens significantly enhances the performance of the vanilla Show-o model by an average of 8.9%, while utilizing fewer tokens. Notably, when compared with unified models with a much larger number of parameters, our model also achieves decent performance over all understanding tasks while keeping a smaller training sample($\sim$10 v.s. $\sim$1,000). Given such promising results, we envision UniCTokens as a potential next-generation paradigm for unified personalized understanding.
@@
Personalized Knowledge-Driven Generation.

### v3

@@ Abstract
Personalized models have demonstrated remarkable success in understanding and generating concepts provided by users. However, existing methods use separate concept tokens for understanding and generation, treating these tasks in isolation. This may result in limitations for generating images with complex prompts. For example, given the concept $\langle bo\rangle$, generating $\langle bo\rangle$ wearing its hat" without additional textual descriptions of its hat. We call this kind of generation personalized attribute-reasoning generation. To address the limitation, we present UniCTokens, a novel framework that effectively integrates personalized information into a unified vision language model (VLM) for understanding and generation. UniCTokens trains a set of unified concept tokens to leverage complementary semantics, boosting two personalized tasks. Moreover, we propose a progressive training strategy with three stages: understanding warm-up, bootstrapping generation from understanding, and deepening understanding from generation to enhance mutual benefits between both tasks. To quantitatively evaluate the unified VLM personalization, we present UnifyBench, the first benchmark for assessing concept understanding, concept generation, and attribute-reasoning generation. Experimental results on UnifyBench indicate that UniCTokens shows competitive performance compared to leading methods in concept understanding, concept generation, and achieving state-of-the-art results in personalized attribute-reasoning generation. Our research demonstrates that enhanced understanding improves generation, and the generation process can yield valuable insights into understanding. Our code and dataset will be released at: https://github.com/arctanxarc/UniCTokens.
Figure 1: The capability overview of UniCTokens. UniCTokens achieves personalized understanding and generation of a unified VLM using user-provided concept images and texts. This is accomplished by fine-tuning a set of unified concept tokens, which harness the mutual benefits of understanding and generation. Notably, UniCTokens supports complex personalized attribute-reasoning generation, which has never been achieved by previous methods.
@@ 1 Introduction
Personalized tasks focus on understanding and generating information that users provide. Recently, there has been growing interest in personalizing understanding models, especially in transforming general-purpose models into daily life assistants talati2024ai ; rahimi2025user ; rahimi2025reasoning . As a result, significant effort has been devoted to improving the personalization capabilities of Large Language Models (LLMs) achiam2023gpt ; bai2023qwen ; grattafiori2024llama and Vision-Language Models (VLMs) liu2023visual ; li2024llava ; li2024llava_in ; bai2023qwen ; zhang2024llama ; zhang2024mathverse ; hong2025worldsense . In terms of generation, with the rapid advancement of text-to-image (T2I) models croitoru2023diffusion ; moser2024diffusion ; guo2025can ; jiang2025t2i ; guo2023point ; li2024zone , personalization techniques can generate highly realistic and diverse images based on user-specified concepts. They mostly employ paradigms such as test-time fine-tuning ruiz2023dreambooth ; zhang2023personalize and retrieval augmentation hao2024remember . LLMs, VLMs, and T2I models have shown remarkable personalization performance in their respective task domains.
Despite significant advancements in personalized generation and understanding, current methods often treat these as independent tasks, failing to effectively leverage complementary semantics nguyen2025yochameleon . Personalized understanding utilizes vision-language models (VLMs) alaluf2024myvlm ; an2024mc ; pi2024personalized , while personalized generation primarily employ diffusion models gal2022image ; ruiz2023dreambooth ; kumari2023multi . This leads to a lack of a unified personalized model for users to perform both understanding and generation. Meanwhile, personalized tasks are complex and conceptually diverse in reality, as shown in Fig. 1. For instance, when training data includes only the concept $\langle bo\rangle$, diffusion models would struggle to generate suitable images of $\langle bo\rangle$ wearing its hat", if additional textual descriptions of its hat are provided. Additionally, concepts often include subtle visual features that are critical for identifying them. Traditional understanding models typically prioritize high-level semantic information over these subtle yet critical features wu2024janus ; chen2025janus . Ignoring the mutual semantics of the two tasks makes current methods insufficient for fully understanding and efficiently generating concepts provided by users.
Although a recent work Yo’Chameleon nguyen2025yochameleon achieves understanding and generation upon a unified VLM team2024chameleon , it still presents several challenges in personalization:
@@
  • Second, Yo’Chameleon assesses personalized understanding and generation separately, without quantifying how understanding facilitates generation, referred as attribute-reasoning generation.
@@
To better assess the personalization capabilities of the unified model, we introduce a new benchmark, UnifyBench, aimed at evaluating models’ abilities in concept understanding, concept generation, and attribute-reasoning generation. When we assess our UniCTokens using UnifyBench, we consistently achieve competitive or better results than all other personalization methods. Our analysis indicates that a better understanding enhances generation, while the generation process can also provide information that supports understanding, providing valuable insights for the development of general unified models. We believe that UnifyBench will serve as a strong foundation for future research in unified model personalization. To sum up, our contributions can be concluded as:
@@
  • We construct UnifyBench, a benchmark to evaluate concept understanding, concept generation, and attribute-reasoning generation of a personalized unified model.
  • We conduct extensive experiments on UnifyBench. UniCTokens demonstrates competitive performance compared to leading methods in concept understanding and generation, achieving state-of-the-art results in attribute-reasoning generation.
@@ 2 Related Work
Model personalization mainly involves integrating concept-related information into the outputs of the model. Recent text-to-image methods depend on recontextualizing text conditions. Text inversion gal2022image utilizes soft prompt tuning for special token adjustments, while Dreambooth ruiz2023dreambooth modifies the entire network weights to ensure subject fidelity. Additionally,  kumari2023multi ; wang2024instantid ; ye2023ip ; zeng2023ipdreamer ; li2024uvidm inject concept-related information through varied training strategies. Large Language Models and Vision-Language Models have also witnessed trends in personalization.  tan2024personalized employs a patch-based LoRA for character, while pi2024personalized utilizes a dual-tower architecture for user-aware outputs. Furthermore,  alaluf2024myvlm ; an2024mc ; hao2024remember achieve personalized responses in multimodal scenarios through fine-tuning or retrieval-augmented generation, linking user information with content in images. Yo’Chameleon nguyen2025yochameleon is the first attempt at unified personalized models. However, its separate training strategy limits cross-task information sharing. We train unified concept tokens to enhance information transfer between understanding and generation tasks.
@@ 3.1 Unified VLM Personalization with Unified Concept Tokens
To model the complexity of real-world personalization, we train unified concept tokens rather than using separate tokens for understanding and generation as Yo’Chameleon nguyen2025yochameleon . Given a target concept $C$, users provide personalized inputs comprising: the name of $C$, a set of images $\{I^{i}\}_{i=1}^{n}$ (typically 3 to 10) that exclusively depict $C$, and a set of textual descriptions $\{T^{i}\}_{i=1}^{m}$ containing additional attributes of $C$ that are not visually inferable (e.g., “$C$’s favorite hat is red”).
The goal is to train unified concept tokens $M$ that can: (1) concept understanding: generate personalized textual responses $P_{\text{text}}$ related to $C$ (e.g., answering “What is $C$ doing in the photo?”), (2) concept generation: synthesize conditional images $P_{\text{img}}$ of $C$ under various prompts, and (3) attribute-reasoning generation: produce personalized attribute-reasoning images $P_{\text{pimg}}$ that integrate the textual information. As shown in Figure 1, the commonality between Object 2 and Object 3 lies in their classification as image generation tasks that necessitate the production of high-quality personalized concepts. The distinction between them is that the prompt for Object 2 contains only the personalized concept, whereas the prompt for Object 3 additionally incorporates certain information that is only within the understanding data (e.g., "[object Object] has a red hat" in understanding data, "a photo of [object Object] wearing its hat" for attribute-reasoning generation). Failure to leverage this information would preclude the generation of the hat in an appropriate color. Our objective is to utilize this task to measure the extent of information transfer across tasks. Detailed task and metric settings can be found in the Appendix. This process can be expressed in a formula as follows:
@@ 3.2 Stage-1: Personalized Understanding Warm-up
Soft prompt tuning has proven effective in integrating new concepts for both personalized understanding and generation tasks nguyen2024yo ; gal2022image . Moreover, learnable prompts are often utilized as conduits for information transfer across task lee2024bayesian ; vu2021spot , model su2021transferability ; belanec2024task , and modality wu2024mixture ; yang2024soft ; xiao2025dynaprompt . Based on this paradigm, we represent the concept $C$ as a prompt containing learnable tokens for unified models:
@@
where $\langle\text{sks}\rangle$ is a learnable unique identifier for this new concept and $\langle\text{tokens}\_{\text{s}}\rangle$ is shared tokens $\langle\text{token}_{\text{s}_{1}}\rangle\dots\langle\text{token}_{\text{s}_{K}}\rangle$ encode semantic attributes specific to that concept. This personalized prompt serves as the system prompt during training. After understanding task training, $\langle\text{tokens}\_{\text{s}}\rangle$ encapsulates various characteristics of the concept (e.g., height, hairstyle, shape; see Fig. 2(2)).
To enable effective cross-task information transfer in subsequent stages, warm-up training is essential. To stabilize the training process, we adopt the token initialization strategy proposed in MC-LLaVA an2024mc . During training, instruction tuning is employed to optimize the initial tokens. The training dataset contains three types of Visual Question Answering (VQA) samples: (1) positive recognition VQA pairs, (2) random recognition VQA pairs, and (3) conversational VQA pairs. Detailed descriptions of the data construction process can be found in MC-LLaVA an2024mc Notably, user-provided textual information $\{T^{i}\}_{i=1}^{m}$ is transformed by LLMs into a set of text-only QA pairs, which are also incorporated into training. Implementation details are provided in the Appendix.
@@ 3.3 Stage-2: Bootstrap Generative Learning from Understanding
Training solely on understanding tasks does not equip the unified model to directly generate images containing $C$. Existing personalized unified models nguyen2025yochameleon require a substantial number of samples ($\sim$1,000) to train a single concept, which is impractical for real-world applications. For humans, a preliminary understanding of concepts facilitates artistic creation. The more thoroughly humans comprehend a concept, the more accurately they can replicate it in their artwork. Thus, we aim to explore the potential of leveraging understanding information to facilitate generation.
Directly training generation tasks on shared tokens presents a straightforward strategy. However, this direct training approach results in the model losing its general conditioning capability on $C$ and significantly diminishes performance on understanding tasks. We posit that this discrepancy is due to variations in task distributions, which encompass specific information that cannot be directly shared. Thus, we integrate new tokens to enhance model training tailored for text-to-image (T2I). Inspired by DreamBooth ruiz2023dreambooth , the prompt for Stage 2 of T2I training is formalized as follows:
$\text{A photo of}\ \langle\text{tokens}\_{\text{s}}\rangle\ \langle\text{tokens}\_{\text{g}}\rangle\ \langle\text{sks}\rangle.$ (3)
@@ 3.4 Stage-3: Deepen Understanding Representation from Generation
Existing literature wu2024janus shows that understanding models focus primarily on high-level semantic information, while the generation model is more responsive to low-level features. Experimental findings indicate that the training in Stage 1 may not effectively address tasks related to low-level characteristics, such as fine-grained recognition. After completing the generation training, model is capable of producing visually similar images, which may have the potential to enhance its understanding capabilities through effective utilization of information derived from generation process.
Show-o xie2024show utilizes a decode methodology following MaskGIT chang2022maskgit , removing masked tokens to derive the final image. To generate an image $x_{T}$, the process can be formularized as follows:
@@
During Stage 3, understanding samples are exclusively constructed from recognition VQA pairs to further strengthen the model’s identification capability. In parallel, the T2I training data continues to update the concept identifier $\langle\text{sks}\rangle$. At this stage, the only learnable tokens are $\langle\text{sks}\rangle$ and $\langle\text{tokens}\_{\text{u}}\rangle$.
@@ 4.1 Experiment Setup
Implement Details. We set the number of learnable tokens as $K=16$, $M=8$, and $N=8$ respectively. All training stages are optimized using AdamW and each stage is trained for 20 epochs. The batch size is set to 4 for understanding tasks in stage 1, and 1 for both stage 2 and stage 3, as well as for T2I generation. All experiments are conducted on A800 GPUs. For the backbone, we adopt Show-o512x512 xie2024show as the base model. More training details can be found in the Appendix.
Dataset. We collect Unifybench of 20 concepts: Person (10), Pets (5), Objects (5). Each concept is associated with 10~15 images for training and testing. In addition, each concept is annotated with 1~2 extra attributes that are not visually inferable from training images (e.g., “this person owns a dog”). To comprehensively evaluate both understanding and T2I capabilities, We design specific test data for each task. This includes standard understanding QA pairs and generation prompts, as well as personalized attribute-reasoning queries. Please refer to Appendix for more details on our dataset.
Baselines. Our approach is primarily evaluated against four distinct categories of methods. The most direct comparison utilizes a unified base model incorporating personalized text and image prompts. Show-o’s inability to support interleaved image prompts precluded this comparison. The textual prompts are derived from the captions generated by GPT-4o for each concept and subsequently summarized to meet the required token length. Another significant baseline for comparison is the recent unified customized model, Yo’chameleon nguyen2025yochameleon . We retrain the model according to the original paper, utilizing 1,000 images per concept and a 7B base model. Additionally, we evaluate the current GPT-4o on our benchmark to serve as an upper bound. Lastly, we conduct comparisons between models focused solely on understanding or generation. More details are provided in the Appendix.
Metrics. For personalized attribute-reasoning image generation, we use VLMs to score the generated images (from 0 to 1) based on their consistency with the extra attributes of concepts embedded in the prompt. Metrics for separate personalized understanding and generation are detailed in Appendix. Final results are obtained by averaging the scores across all concepts for each evaluation metric.
@@ 4.2 Our Unified Personalized Benchmark (UnifyBench)
Type Methods Model Size Token Training Images Personalized Understanding Personalized Generation PARG Rec. VQA QA Pure Gen. People Gen. Weight BLEU GPT BLEU GPT CLIP-I CLIP-T DINO Face-Simi Score CLIP-I Upper Bound GPT-4o+TP 200B $\sim$100 - 0.742 0.473 0.676 0.610 0.685 0.689 0.301 0.626 0.198 0.780 0.690 GPT-4o+IP 200B $\sim$1,000 - 0.773 0.543 0.685 0.589 0.652 0.794 0.310 0.722 0.559 0.782 0.792 Real Images - - - - - - - - 0.832 - 0.728 0.739 - 0.832 Und. Only Yo’LLaVA 13B 16 $\sim$100 0.919 0.609 0.629 0.612 0.593 - - - - - - MC-LLaVA 13B 16 $\sim$10 0.924 0.628 0.637 0.601 0.583 - - - - - - RAP-MLLM 13B $\sim$1,000 - 0.940 0.616 0.616 0.712 0.722 - - - - - - Qwen2.5-VL + TP 3B $\sim$100 - 0.660 0.407 0.727 0.574 0.774 - - - - - - Yo’LLaVA(Phi-1.5) 1.3B 16 $\sim$ 100 0.765 0.488 0.497 0.510 0.494 - - - - - - Gen. Only Text inversion 1.0B - $\sim$10 - - - - - 0.630 0.247 0.569 0.371 0.070 0.628 DreamBooth (SD) 1.0B - $\sim$10 - - - - - 0.649 0.281 0.591 0.436 0.071 0.650 Unified Model Chamaleon+TP 7B $\sim$100 - 0.690 0.413 0.488 0.509 0.564 0.547 0.176 0.509 0.011 0.329 0.549 Chameleon+IP 7B $\sim$1,000 - 0.493 0.445 0.498 0.407 0.535 0.523 0.160 0.469 0.066 0.299 0.499 Show-o+TP 1.3B $\sim$100 - 0.566 0.461 0.409 0.504 0.579 0.664 0.264 0.553 0.048 0.770 0.660 Yo’Chameleon 7B 32 $\sim$1,000 0.764 0.474 0.507 0.510 0.581 0.697 0.236 0.590 0.224 0.266 0.698 Ours 1.3B 32 $\sim$10 0.790 0.503 0.523 0.544 0.603 0.750 0.282 0.646 0.334 0.359 0.749
Table 1: Quantitative Comparison on UnifyBench. TP = Text Prompt. IP = Image Prompt. PARG = Personalized Attribute-Reasoning Generation. The best and second best are highlighted.
@@
This task requires responding to queries with concept identifiers and images. Following alaluf2024myvlm ; nguyen2024yo ; an2024mc , we conduct experiments on personalized recognition, VQA, and QA tasks. As demonstrated in Tab. 1, our proposed UniCTokens significantly enhances the performance of the vanilla Show-o model by an average of 8.9%, while utilizing fewer tokens. Notably, when compared with unified models with a much larger number of parameters, our model also achieves decent performance over all understanding tasks while keeping a smaller training sample($\sim$10 v.s. $\sim$1,000). Given such promising results, we envision UniCTokens as a potential next-generation paradigm for unified personalized understanding.
@@
Personalized Attribute-Reasoning Generation.